# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v2

@@ VII Conclusion and Future Work
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal viewpoint based on the task goal, then selects the manipulation action from the updated observation. To enhance performance and training stability, we integrate voxel alignment for spatial consistency, demonstration augmentation for enhanced learning, and auxiliary rewards for real-time training guidance. These strategies enable the model to handle occlusions and improve sensor-motor coordination. Experimental results show that our model significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment. However, complex occlusions and dynamic environments remain challenging. Future work could explore real-time visual feedback to better adapt to environmental changes.

### v3

@@ VII Conclusion and Future Work
In this paper, we investigate the problem of robotic manipulation under limited observation and present a task-driven asynchronous active vision-action model. Our model first infers the optimal viewpoint to actively observe the scene, then selects the manipulation action from the updated observation. Our model was trained and evaluated on eight RLBench tasks with varying levels of occlusion, demonstrating that it significantly outperforms passive observation-based models by actively refining viewpoints and learning sensor-motor coordination through interaction with the environment. However, one limitation is its reliance on a small number of demonstration samples to provide early training signals. Future work could explore strategies for zero-shot training and extend the approach to longer-horizon tasks.